a particle of mass M carrying charge +q1 is revolving around a fixed charge -q2 in circular path of radius r.calculate the period of revolution.

Grade:11

2 Answers

umesh chandrakant shah
13 years ago

Post your question at right forum.

 

gOlU g3n|[0]uS
13 years ago

Fe=Centripital force

k*q1*q2/r2 = Mrw2  [w=omega]

w2= k*q1*q2/Mr3

(2Π/T )2= k*q1*q2/Mr3

T2 = 4Π2 Mr3/k*q1*q2

T = (4Π2 Mr3/k*q1*q2)1/2